# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 2.7. Temporal valued structures
The following theorem is a special case of [15, Theorem 3.4].
Theorem 2.25.
@@
We will often use the following notation in the proofs.
Definition 2.26 ($E_{t},N_{t},O_{t}$).
@@
$E_{t}$ :=$:=\{(p,q)\in[k]^{2}\mid t_{p}=t_{q}\},$
$N_{t}$ :=$:=\{(p,q)\in[k]^{2}\mid t_{p}\neq t_{q}\},\text{ and}$
$O_{t}$ :=$:=\{(p,q)\in[k]^{2}\mid t_{p}<t_{q}\},$
where $<$ is the natural order over $\mathbb{Q}$.
@@
3. General Facts
In this section we formulate and prove some relatively easy, but very general and useful facts for analysing the computational complexity of VCSPs.
Lemma 3.1.
@@ 3. General Facts
Proof.
Suppose that there exists $b\in A$ such that the unary operation $c$ defined by $c(a)=b$ for all $a\in A$ is a fractional polymorphism of ${\mathfrak{A}}$. Then for every $R\in\tau$ of arity $k$ and $t\in A^{k}$, we have $R(b,\dots,b)=R(c(t))\leq R(t)$. Let $\phi(x_{1},\dots,x_{n})=\sum_{i}\phi_{i}$ be an instance of $\operatorname{VCSP}({\mathfrak{A}})$, where each $\phi_{i}$ is an atomic $\tau$-expression. Then the minimum of $\phi$ equals $\sum_{i}\phi_{i}(b,\dots,b)$ and hence $\operatorname{VCSP}(\Gamma)$ is in P. ∎
@@
Proposition 3.2.
Let ${\mathfrak{A}}$ be a valued structure and let ${\mathfrak{B}}$ be a relational $\tau$-structure on countable domains $A$ and $B$, respectively. Then ${\mathfrak{A}}$ pp-constructs ${\mathfrak{B}}$ if and only if $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ pp-constructs ${\mathfrak{B}}$.
Proof.
Clearly, whenever $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ pp-constructs ${\mathfrak{B}}$, then ${\mathfrak{A}}$ pp-constructs ${\mathfrak{B}}$. Suppose that ${\mathfrak{A}}$ pp-constructs ${\mathfrak{B}}$. Then there exists $d\in\mathbb{N}$ and a pp-power ${\mathfrak{C}}$ on the domain $C=A^{d}$ of ${\mathfrak{A}}$ which is fractionally homomorphically equivalent to ${\mathfrak{B}}$. We claim that $\operatorname{Feas}({\mathfrak{C}})$ is fractionally homomorphically equivalent to ${\mathfrak{B}}$ as well, witnessed by the same fractional homomorphisms.
@@
is finite. Since $R^{{\mathfrak{B}}}$ attains only values 00 and $\infty$, it follows that $E_{\omega_{1}}[f\mapsto R^{{\mathfrak{B}}}(f(t))]=0$. Therefore (4) holds. Since $R$ and $t$ were arbitrary, it follows that $\omega_{1}$ is a fractional homomorphism from $\operatorname{Feas}({\mathfrak{C}})$ to ${\mathfrak{B}}$. The proof that $\omega_{2}$ is a fractional homomorphism from ${\mathfrak{B}}$ to $\operatorname{Feas}({\mathfrak{C}})$ is similar.
Note that $\operatorname{Feas}({\mathfrak{C}})$ is a pp-power of $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$: every relation $R$ of ${\mathfrak{C}}$ of arity $k$ lies in $\langle{\mathfrak{A}}\rangle$ when viewed as a relation of arity $dk$ and therefore $\operatorname{Feas}(R)\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$. Hence, $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ pp-constructs ${\mathfrak{B}}$ as we wanted to prove. ∎
@@
Conversely, suppose that ${\mathfrak{A}}$ is essentially crisp. Let $R$ be a valued relation of arity $k$ in ${\mathfrak{A}}$ and let $s,t\in A^{k}$. If $R(s)=\infty$ or $R(t)=\infty$ then $R(\pi^{2}_{1}(s,t))\leq\frac{R(s)+R(t)}{2}=\infty$. Otherwise, $R(s)=R(t)$ and the inequality holds trivially. ∎
4. Equality VCSPs
@@
$\operatorname{Dis}$ :=$:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x=y\neq z)\vee(x\neq y=z)\}.$ (5)
It is known that $(\mathbb{Q};\operatorname{Dis})$ pp-constructs $K_{3}$, see, e.g., Theorem 7.4.1 and Corollary 6.1.23 in [3]. Let $\operatorname{const}\colon{\mathbb{Q}}\to{\mathbb{Q}}$ be the constant zero operation, given by $\operatorname{const}(x):=0$ for all $x\in{\mathbb{Q}}$. Let $\operatorname{inj}\colon\mathbb{Q}^{2}\to\mathbb{Q}$ be injective. A tuple is called injective if it has pairwise distinct entries.
Theorem 4.1 ([3, 8]).
@@
  • $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$ or $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$. In this case, for every reduct ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{A}}$ with a finite signature, $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is in P.
  • The relation $\operatorname{Dis}$ has a primitive positive definition in ${\mathfrak{A}}$. In this case, ${\mathfrak{A}}$ pp-constructs $K_{3}$ and ${\mathfrak{A}}$ has a reduct ${\mathfrak{A}}^{\prime}$ with a finite signature such that $\operatorname{CSP}({\mathfrak{A}})$ is NP-complete.
We prove the following general lemma that assumes only $\operatorname{Aut}(\mathbb{Q};<)\subseteq\operatorname{Aut}({\mathfrak{A}})$ to avoid repeating the proof in Section 5. The case $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Sym}(\mathbb{Q})$ is a special case.
Lemma 4.2.
@@
Let $a\in A$ and note that, by 2.27, value $\psi(a,a)$ does not depend on the choice of $a$. By our choice of $i$, $\psi(a,a)>\psi(t_{i-1},t_{i})$; otherwise, there are $a_{1},\dots,a_{i-2}\in A$ such that $\psi_{i}(a_{1},\dots,a_{i-2},a,a)\leq\psi_{i}(t_{1},\dots,t_{i})$, in contradiction to the choice of $(t,i)$ such that $i$ is minimal. We distinguish three cases (recall that $t_{i-1}\neq t_{i}$):
  1. (1)
@@
Note that for $a,b\in A$ such that $a<b$, the values $\psi(a,b)$ and $\psi(b,a)$ do not depend on the choice of $a$, $b$. In case (1), $\operatorname{Opt}(\psi)$ expresses $(\neq)_{0}^{\infty}$. In case (2), $\operatorname{Opt}(\psi)$ expresses $(<)_{0}^{\infty}$. Finally, in case (3) $\operatorname{Opt}(\psi)$ expresses $(>)_{0}^{\infty}$, which expresses $(<)_{0}^{\infty}$ by exchanging the input variables.
The last statement follows from the fact that $\operatorname{Sym}(\mathbb{Q})$ does not preserve $(<)_{0}^{\infty}$. ∎
Lemma 4.3 (see, e.g., [15, Example 4.9]).
Let ${\mathfrak{A}}$ be a valued structure such that $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Sym}(\mathbb{Q})$. If $\langle{\mathfrak{A}}\rangle$ contains $(=)^{1}_{0}$ and $(\neq)_{0}^{\infty}$, then $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$. In particular, ${\mathfrak{A}}$ pp-constructs $K_{3}$, and, if the signature of ${\mathfrak{A}}$ is finite, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
Lemma 4.4.
@@
We first prove that $\psi(g(x))$ is finite. Let $u^{1},\dots,u^{n}$ be an enumeration of representatives of all orbits of $k$-tuples such that $\psi(u^{i})<\infty$. If for some distinct $p,q\in\{1,\dots,k\}$ we have $(u^{i})_{p}=(u^{i})_{q}$ for all $i\in\{1,\dots,n\}$, then the algorithm would have replaced all occurrences of $x_{p}$ by $x_{q}$ or vice versa. So for all distinct $p,q\in\{1,\dots,k\}$ there exists $i\in\{1,\dots,n\}$ such that $(u^{i})_{p}\neq(u^{i})_{q}$. Therefore, since $\operatorname{inj}$ is injective, the tuple $\operatorname{inj}(u^{1},\operatorname{inj}(u^{2},\dots,\operatorname{inj}(u^{n-1},u^{n})\dots))$ lies in the same orbit as $g(x)$. Since $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, we have $\psi(g(x))<\infty$.
@@
  • $\operatorname{inj}(g(x),f(x))$ lies in the same orbit of $\operatorname{Aut}({\mathfrak{A}})$ as $g(x)$, and thus $\psi(\operatorname{inj}(g(x),f(x))=\psi(g(x))$.
Combining, we obtain that $\psi(g(x))\leq\psi(f(x))$. It follows that $g$ minimises $\phi^{\prime}$. Recall that the minimum of $\phi$ and $\phi^{\prime}$ are equal. Therefore, the algorithm accepts if the evaluation of $\phi^{\prime}$ under $g$ is at most $u$ and rejects otherwise. Since checking whether a summand forces an equality can be done in constant time, and there is a linear number of variables, the propagation of forced equalities can be done in polynomial time. It follows that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. ∎
Lemma 4.5.
@@
Note that, in particular, $R(s),R(t)<\infty$. Suppose first that $\operatorname{inj}(s,t)=\infty$. In this case, the inequality above implies that $\operatorname{Feas}(R)$ is not improved by $\operatorname{inj}$. It follows that $\operatorname{Pol}(\mathbb{Q};\operatorname{Feas}(R),(\neq)_{0}^{\infty})$ contains neither $\operatorname{const}$ nor $\operatorname{inj}$. Hence, by Theorem 4.1, $(\mathbb{Q};\operatorname{Feas}(R),(\neq)_{0}^{\infty})$ primitively positively defines $\operatorname{Dis}$ and thus $\operatorname{Dis}\in\langle({\mathfrak{A}},(\neq)_{0}^{\infty}\rangle$. We may therefore assume in the rest of the proof that $R(\operatorname{inj}(s,t))<\infty$.
@@
Out of all pairs $(s,t)\in(A^{k})^{2}$ such that $s$ and $\operatorname{inj}(s,t)$ lie in the same orbit and $R(t)<R(s)$, we choose $(s,t)$ such that $t_{p}\neq t_{q}$ holds for as many pairs $(p,q)$ as possible. Note that $s$ and $t$ cannot lie in the same orbit, and by the injectivity of $\operatorname{inj}$ there exist $i,j\in\{1,\dots,k\}$ such that $t_{i}=t_{j}$ and $s_{i}\neq s_{j}$. Note that since $s$ and $\operatorname{inj}(s,t)$ lie in the same orbit, we have $E_{s}\subseteq E_{t}$ and $N_{t}\subseteq N_{s}$ (recall Definition 2.26). For the sake of the notation, assume that $i=1$ and $j=2$. Consider the expression
@@
By (iii), $\operatorname{inj}(s,u)$ lies in the same orbit as $s$. By (i), we get that $R(u)\leq R(t)<R(s)$. By (ii) and (iv), $u$ satisfies $u_{p}\neq u_{q}$ for more pairs $(p,q)$ than $t$, which contradicts our choice of $t$. Therefore, $m<\ell$. It follows that $\phi(x_{1},x_{2})$ is equivalent to $(=)_{m}^{\ell}$ with $m<\ell$. Recall that $(=)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$ by definition and therefore $(=)_{m}^{\ell}\in\langle{\mathfrak{A}},(\neq)_{0}^{\infty}\rangle$. Note that $\ell\leq R(s)<\infty$. Hence, shifting $\phi$ by $-m$ and scaling it by $1/(\ell-m)$ shows that $(=)^{1}_{0}\in\langle{\mathfrak{A}},(\neq)_{0}^{\infty}\rangle$, as we wanted to prove. ∎
Theorem 4.6.
@@
  1. (1)
    $(\neq)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$ and $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$. In this case, ${\mathfrak{A}}$ pp-constructs $K_{3}$, and $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
  2. (2)
@@
Proof.
If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then Lemma 3.1 implies that $\operatorname{VCSP}(\Gamma)$ is in P. If $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$, then ${\mathfrak{A}}$ can express $(\neq)_{0}^{\infty}$ by Lemma 4.2.
If $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, then Lemma 4.4 implies that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. If $\operatorname{inj}\notin\operatorname{fPol}({\mathfrak{A}})$, then ${\mathfrak{A}}$ can express $\operatorname{Dis}$ or $(=)_{0}^{1}$ by Lemma 4.5. If $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$, then the statement follows from Theorem 4.1. If $(=)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$, then we obtain that $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$ by Lemma 4.3 and again the statement follows. Note that neither $\operatorname{const}$ nor $\operatorname{inj}$ improves $\operatorname{Dis}$. Therefore, the two cases in the statement of the theorem are disjoint. ∎
Remark 4.7.
If ${\mathfrak{A}}$ is as in Theorem 4.6, then ${\mathfrak{A}}$ has a first-order definition in $(\mathbb{Q};=)$ as introduced in Section 2.3, where the defining formulas can be chosen as disjunctions of conjunctions of atomic formulas over $(\mathbb{Q};=)$. If ${\mathfrak{A}}$ is given by such a first-order definition over $(\mathbb{Q};=)$, then it is decidable which of the conditions (1) and (2) in Theorem 4.6 applies: we can just test whether all valued relations of ${\mathfrak{A}}$ are improved by $\operatorname{const}$, and we can test whether all of them are improved by $\operatorname{inj}$.
Remark 4.8.
The complexity classification of equality minCSPs from [35], which can be viewed as VCSPs of valued structures where each relation attains only values 00 and $1$, can be obtained as a special case of Theorem 4.6. Suppose that ${\mathfrak{A}}$ is such a valued structure. If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then ${\mathfrak{A}}$ is constant (in the terminology of [35]) and $\operatorname{VCSP}({\mathfrak{A}})$ is in P. If $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, then it is immediate that ${\mathfrak{A}}$ is Horn (in the terminology of [35]) and even strictly negative: otherwise, by [35, Lemma 16] we have $(=)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. But this is in contradiction to the assumption that $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, since $\operatorname{inj}$ applied to a pair of equal elements and pair of distinct elements yields a pair of distinct elements, increasing the cost to $1$ compared to the average cost $1/2$ of the input tuples. Otherwise, it follows from Theorem 4.6 that $\operatorname{VCSP}({\mathfrak{A}})$ is NP-hard.
5. Temporal VCSPs

### v2

@@ 2.7 Temporal valued structures
The structure $(\mathbb{Q};<)$ is known to be finitely bounded and homogeneous, which yields the following theorem as a special case of [16, Theorem 3.4].
Theorem 21
@@
To increase readability of some of the proofs, we we will often use the following notation.
Definition 22 ($E_{t},N_{t},O_{t}$)
@@
$E_{t}$ $:=\{(p,q)\in[k]^{2}\mid t_{p}=t_{q}\},$
$N_{t}$ $:=\{(p,q)\in[k]^{2}\mid t_{p}\neq t_{q}\},\text{ and}$
$O_{t}$ $:=\{(p,q)\in[k]^{2}\mid t_{p}<t_{q}\},$
where << is the natural order over $\mathbb{Q}$.
@@
3 General facts
In this section we prove some relatively easy, but very general and useful facts for analysing the computational complexity of VCSPs.
Lemma 24
@@ 3 General facts
Proof 2
Suppose that there exists $b\in A$ such that the unary operation $c$ defined by $c(a)=b$ for all $a\in A$ is a fractional polymorphism of $\mathfrak{A}$}. Then for every $R\in\tau$ of arity $k$ and $t\in A^{k}$, we have $R(b,\dots,b)=R(c(t))\leq R(t)$. Let $\phi(x_{1},\dots,x_{n})=\sum_{i}\phi_{i}$ be an instance of $\operatorname{VCSP}({\mathfrak{A}})$, where each $\phi_{i}$ is an atomic $\tau$-expression. Then the minimum of $\phi$ equals $\sum_{i}\phi_{i}(b,\dots,b)$ and hence $\operatorname{VCSP}({\mathfrak{A}})$ is in P. $\qed$
@@
Proposition 25
Let $\mathfrak{A}$} be a valued structure and let $\mathfrak{B}$} be a relational $\tau$-structure on countable domains $A$ and $B$, respectively. Then $\mathfrak{A}$} $p$-constructs $\mathfrak{B}$} if and only if $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ $p$-constructs $\mathfrak{B}$}.
Proof 3
Clearly, if $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ $p$-constructs $\mathfrak{B}$}, then $\mathfrak{A}$} $p$-constructs $\mathfrak{B}$}. Suppose that $\mathfrak{A}$} $p$-constructs $\mathfrak{B}$}. Then there exists $d\in\mathbb{N}$ and a pp-power $\mathfrak{C}$} on the domain $C=A^{d}$ of $\mathfrak{A}$} which is fractionally homomorphically equivalent to $\mathfrak{B}$}. We claim that $\operatorname{Feas}({\mathfrak{C}})$ is fractionally homomorphically equivalent to $\mathfrak{B}$} as well, witnessed by the same fractional homomorphisms.
@@
is finite. Since $R^{{\mathfrak{B}}}$ attains only values 0 and $\infty$, it follows that $E_{\omega_{1}}[f\mapsto R^{{\mathfrak{B}}}(f(t))]=0$. Therefore (3) holds. Since $R$ and $t$ were arbitrary, it follows that $\omega_{1}$ is a fractional homomorphism from $\operatorname{Feas}({\mathfrak{C}})$ to $\mathfrak{B}$}. The proof that $\omega_{2}$ is a fractional homomorphism from $\mathfrak{B}$} to $\operatorname{Feas}({\mathfrak{C}})$ is similar.
Note that $\operatorname{Feas}({\mathfrak{C}})$ is a pp-power of $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$: every relation $R$ of $\mathfrak{C}$} of arity $k$ lies in $\langle{\mathfrak{A}}\rangle$ when viewed as a relation of arity $dk$ and therefore $\operatorname{Feas}(R)\in\langle{\mathfrak{A}}\rangle_{0}^{\infty}$. Hence, $(A;\langle{\mathfrak{A}}\rangle_{0}^{\infty})$ $p$-constructs $\mathfrak{B}$} as we wanted to prove. $\qed$
@@
Conversely, suppose that $\mathfrak{A}$} is essentially crisp. Let $R$ be a valued relation of arity $k$ in $\mathfrak{A}$} and let $s,t\in A^{k}$. If $R(s)=\infty$ or $R(t)=\infty$ then $R(\pi^{2}_{1}(s,t))\leq\frac{R(s)+R(t)}{2}=\infty$. Otherwise, $R(s)=R(t)$ and the inequality holds trivially. $\qed$
4 Equality VCSPs
@@
$\operatorname{Dis}$ $:=\{(x,y,z)\in\mathbb{Q}^{3}\mid(x=y\neq z)\vee(x\neq y=z)\}.$ (4)
It is known that $(\mathbb{Q};\operatorname{Dis})$ $p$-constructs $K_{3}$, see, e.g., Theorem 7.4.1 and Corollary 6.1.23 in [17]. Let $\operatorname{const}\colon{\mathbb{Q}}\to{\mathbb{Q}}$ be the constant zero operation, given by $\operatorname{const}(x):=0$ for all $x\in{\mathbb{Q}}$. Let $\operatorname{inj}\colon\mathbb{Q}^{2}\to\mathbb{Q}$ be injective. A tuple is called injective if it has pairwise distinct entries.
Theorem 27 ([17, 7])
@@
  • 1.
    $\operatorname{const}\in\operatorname{Pol}({\mathfrak{A}})$ or $\operatorname{inj}\in\operatorname{Pol}({\mathfrak{A}})$. In this case, for every reduct $\mathfrak{A}$$^{\prime}$ of $\mathfrak{A}$} with a finite signature, $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is in P.
  • 2.
    The relation $\operatorname{Dis}$ has a primitive positive definition in $\mathfrak{A}$}. In this case, $\mathfrak{A}$} $p$-constructs $K_{3}$ and $\mathfrak{A}$} has a reduct $\mathfrak{A}$$^{\prime}$ with a finite signature such that $\operatorname{CSP}({\mathfrak{A}}^{\prime})$ is NP-complete.
The complexity classification of equality VCSPs is a consequence of several lemmas. The statements are inspired by analogous statements for equality CSPs that link the pp-definability of certain relations with the non-existence of certain polymorphisms [7]. Our proofs, however, are quite remarkable, because we do not know yet whether a general preservation theorem holds that would link the pp-expressibility of valued relations with the non-existence of certain fractional polymorphisms; see the discussion in Remark 19. Our first lemma shows a useful consequence of $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$.
Lemma 28
@@
Let $a\in A$ and note that, by 23, value $\psi(a,a)$ does not depend on the choice of $a$. By our choice of $i$, $\psi(a,a)>\psi(t_{i-1},t_{i})$; otherwise, there are $a_{1},\dots,a_{i-2}\in A$ such that $\psi_{i}(a_{1},\dots,a_{i-2},a,a)\leq\psi_{i}(t_{1},\dots,t_{i})$, in contradiction to the choice of $(t,i)$ such that $i$ is minimal. We distinguish three cases (recall that $t_{i-1}\neq t_{i}$):
  1. 1.
@@
Note that for $a,b\in A$ such that $a<b$, the values $\psi(a,b)$ and $\psi(b,a)$ do not depend on the choice of $a$, $b$. In case 1, $\operatorname{Opt}(\psi)$ is equal to $(\neq)_{0}^{\infty}$. In case 2, $\operatorname{Opt}(\psi)$ equals $(<)_{0}^{\infty}$. Finally, in case 3 $\operatorname{Opt}(\psi)$ is equal to $(>)_{0}^{\infty}$, from which we obtain $(<)_{0}^{\infty}$ by exchanging the input variables.
The last statement follows from the fact that $\operatorname{Sym}(\mathbb{Q})$ does not preserve $(<)_{0}^{\infty}$. $\qed$
As a next step, we prove tractability for equality valued structures with an injective fractional polymorphism. The high level idea is that to find a solution to an instance, we collapse variables whose equality is forced and find an injective solution for the remaining variables.
Lemma 29
@@
We first prove that $\psi(g(x))$ is finite. Let $u^{1},\dots,u^{n}$ be an enumeration of representatives of all orbits of $k$-tuples such that $\psi(u^{i})<\infty$ and note that $n\geq 1$, because otherwise the algorithm would have rejected the instance. If for some distinct $p,q\in\{1,\dots,k\}$ we have $(u^{i})_{p}=(u^{i})_{q}$ for all $i\in\{1,\dots,n\}$, then the algorithm would have replaced all occurrences of $x_{p}$ by $x_{q}$ or vice versa. So for all distinct $p,q\in\{1,\dots,k\}$ there exists $i\in\{1,\dots,n\}$ such that $(u^{i})_{p}\neq(u^{i})_{q}$. Therefore, since $\operatorname{inj}$ is injective, the tuple $\operatorname{inj}(u^{1},\operatorname{inj}(u^{2},\dots,\operatorname{inj}(u^{n-1},u^{n})\dots))$ lies in the same orbit as $g(x)$. Since $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, we have $\psi(g(x))<\infty$.
@@
  • 2.
    $\operatorname{inj}(g(x),f(x))$ lies in the same orbit of $\operatorname{Aut}({\mathfrak{A}})$ as $g(x)$, and thus
    $\psi(\operatorname{inj}(g(x),f(x))=\psi(g(x)).$
Combining, we obtain that $\psi(g(x))\leq\psi(f(x))$. It follows that $g$ minimises $\phi^{\prime}$. Recall that the minimum of $\phi$ and $\phi^{\prime}$ are equal. Therefore, the algorithm accepts if the evaluation of $\phi^{\prime}$ under $g$ is at most $u$ and rejects otherwise. Since checking whether a summand forces an equality can be done in constant time, and there is a linear number of variables, the propagation of forced equalities can be done in polynomial time. It follows that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. $\qed$
The last auxiliary statement shows that if $\operatorname{inj}\notin\operatorname{fPol}({\mathfrak{A}})$, then this has strong consequences for expressive power of $\mathfrak{A}$}. The proof idea is to use the existence of tuples that certify that $\operatorname{inj}\notin\operatorname{fPol}({\mathfrak{A}})$ to show that we can express some valued relations.
Lemma 30
@@
Note that, in particular, $R(s),R(t)<\infty$. Suppose first that $R(\operatorname{inj}(s,t))=\infty$. In this case, the inequality above implies that $\operatorname{Feas}(R)$ is not improved by $\operatorname{inj}$. It follows that $\operatorname{Pol}(\mathbb{Q};\operatorname{Feas}(R),(\neq)_{0}^{\infty})$ contains neither $\operatorname{const}$ nor $\operatorname{inj}$. Hence, by Theorem 27, $\operatorname{Dis}$ is primitively positively definable in $(\mathbb{Q};\operatorname{Feas}(R),(\neq)_{0}^{\infty})$ and thus $\operatorname{Dis}\in\langle({\mathfrak{A}},(\neq)_{0}^{\infty}\rangle$. We may therefore assume in the rest of the proof that $R(\operatorname{inj}(s,t))<\infty$.
@@
Out of all pairs $(s,t)\in(A^{k})^{2}$ such that $s$ and $\operatorname{inj}(s,t)$ lie in the same orbit and $R(t)<R(s)$, we choose $(s,t)$ such that $t_{p}\neq t_{q}$ holds for as many pairs $(p,q)$ as possible. Note that $s$ and $t$ cannot lie in the same orbit, and by the injectivity of $\operatorname{inj}$ there exist $i,j\in\{1,\dots,k\}$ such that $t_{i}=t_{j}$ and $s_{i}\neq s_{j}$. Note that since $s$ and $\operatorname{inj}(s,t)$ lie in the same orbit, we have $E_{s}\subseteq E_{t}$ and $N_{t}\subseteq N_{s}$ (recall definition 22). For the sake of the notation, assume that $i=1$ and $j=2$. Consider the pp-expression
@@
By (iii), $\operatorname{inj}(s,u)$ lies in the same orbit as $s$. By (i), we get that $R(u)\leq R(t)<R(s)$. By (ii) and (iv), $u$ satisfies $u_{p}\neq u_{q}$ for more pairs $(p,q)$ than $t$, which contradicts our choice of $t$. Therefore, $m<\ell$. It follows that $\phi(x_{1},x_{2})$ is equivalent to $(=)_{m}^{\ell}$ with $m<\ell$. Recall that $(=)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$ by definition and therefore $(=)_{m}^{\ell}\in\langle{\mathfrak{A}},(\neq)_{0}^{\infty}\rangle$. Note that $\ell\leq R(s)<\infty$. Hence, shifting $\phi$ by m-m and scaling it by $1/(\ell-m)$ shows that $(=)^{1}_{0}\in\langle{\mathfrak{A}},(\neq)_{0}^{\infty}\rangle$, as we wanted to prove. $\qed$
In the proof of the classification theorem for equality VCSPs below we use the following fact.
Lemma 31 (see, e.g., [16, Example 4.9])
Let $\mathfrak{A}$} be a valued structure such that $\operatorname{Aut}({\mathfrak{A}})=\operatorname{Sym}(\mathbb{Q})$. If $\langle{\mathfrak{A}}\rangle$ contains $(=)^{1}_{0}$ and $(\neq)_{0}^{\infty}$, then $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$. In particular, $\mathfrak{A}$} $p$-constructs $K_{3}$, and, if the signature of $\mathfrak{A}$} is finite, $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
Theorem 32
@@
  1. 1.
    $(\neq)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$ and $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$. In this case, $\mathfrak{A}$} $p$-constructs $K_{3}$, and $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete.
  2. 2.
@@
Proof 8
If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then Lemma 24 implies that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. So we suppose that $\operatorname{const}\notin\operatorname{fPol}({\mathfrak{A}})$ in the following; then $(\neq)_{0}^{\infty}\in\langle{\mathfrak{A}}\rangle$ by Lemma 28. If $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, then Lemma 29 implies that $\operatorname{VCSP}({\mathfrak{A}})$ is in P. Therefore we also suppose that $\operatorname{inj}\notin\operatorname{fPol}({\mathfrak{A}})$ in the following. By Lemma 30, we have $(=)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$ or $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$. In fact, Lemma 31 implies that $\operatorname{Dis}\in\langle{\mathfrak{A}}\rangle$ in both cases. Then $\mathfrak{A}$} $p$-constructs $K_{3}$ and $\operatorname{VCSP}({\mathfrak{A}})$ is NP-complete by Theorem 27. Note that neither $\operatorname{const}$ nor $\operatorname{inj}$ improves $\operatorname{Dis}$. Therefore, the two cases in the statement of the theorem are disjoint. $\qed$
Remark 33
The complexity classification of equality minCSPs from [37], which can be viewed as VCSPs of valued structures where each relation attains only values 0 and $1$, can be obtained as a special case of Theorem 32. Suppose that $\mathfrak{A}$} is such a valued structure. If $\operatorname{const}\in\operatorname{fPol}({\mathfrak{A}})$, then $\mathfrak{A}$} is constant (in the terminology of [37]) and $\operatorname{VCSP}({\mathfrak{A}})$ is in P. If $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, then it is immediate that $\mathfrak{A}$} is Horn (in the terminology of [37]) and even strictly negative: otherwise, by [37, Lemma 16] we have $(=)_{0}^{1}\in\langle{\mathfrak{A}}\rangle$. But this is in contradiction to the assumption that $\operatorname{inj}\in\operatorname{fPol}({\mathfrak{A}})$, since $\operatorname{inj}$ applied to a pair of equal elements and pair of distinct elements yields a pair of distinct elements, increasing the cost to $1$ compared to the average cost $1/2$ of the input tuples. Otherwise, it follows from Theorem 32 that $\operatorname{VCSP}({\mathfrak{A}})$ is NP-hard.
5 Temporal VCSPs